OpenStudy (anonymous):

the value of y varies directly with x, and y=4 when x=-10. find y when x= 35.

OpenStudy (victoriasushchik):

y varies directly with x y=kx 10=k(15) k=10/15 y=(10/15)x when x=6 y=(10/15)(6)=60/15=4 x varies inversely as y x=k/y 2=k/8 k=16 x=16/y when y=17 x=16/17